OpenStudy (anonymous):

find angle x in 3 cot x-2 cos x=0

OpenStudy (thomas5267):

\[ \cot(x)=\frac{\cos(x)}{\sin(x)} \] I will then factorise it.

OpenStudy (anonymous):

great!! thanks :)